Ms. Child is considering the purchase of a new food packaging system. The system costs $73,623. Ms. Child plans to borrow one-third of the purchase price from a bank at 4.5% per year compounded annually. The loan will be repaid using equal, annual payments over a 7-year period. The system is expected to last 15 years and have a salvage value of $19,603 at that time. Over the 15 year period, Ms. Child expects to pay $790 per year for maintenance. The system will save $2,444 per year because of efficiencies. Ms. Child uses a MARR of 8% to evaluate investments. What is the equivalent uniform annual worth (EUAW) of this system?

Enter your answer as follows: 1234

Round your answer. Do not use a dollar sign ("$"), any commas (",") or a decimal point (".").

Hint: Identify which information is relevant to answering this question. Draw a CFD.
Answer #1

Laon amount = 73623 / 3 = 24541

Annual loan repayment amount = 24541 * (A/P,4.5%, 7) = 24541 * 0.169701 = 4164.63

Actual investment = 73623 / 3 * 2 = 49082

EUAW = -49082 * (A/P.8%,15) - 4164.63 * (P/A, 8%,7)*(A/P,8%,15) + 19603 * (A/F, 8%,15) + 2444 - 790

= -49082 * 0.116830 - 4164.63 * 5.206370*0.116830 + 19603 *0.036830 + 2444 - 790

= -5891.45 ~-5891 (Nearest Dollar)